Q.

The random variable X follows binomial distribution B(n,p), for which the difference of the mean and the variance is 1. If 2P(X=2)=3P(X=1), then n2P(X>1) is equal to                      [2023]

1 15  
2 11  
3 16  
4 12  

Ans.

(2)

np-npq=1

⇒np(1-q)=1

⇒np2=1                            [∵ p=1-q]      ...(i)

Now, 2P(X=2)=3P(X=1)                           ...(ii)

⇒2·C2n p2qn-2=3·C1n pq n-1                

⇒np-p=3q                              [∵ q=1-p]

⇒np+2p=3                       [from (i) and (ii)]

⇒p=12

From (i), np2=1

⇒n×(12)2=1 ⇒n=4

P(X>1)=1-P(X=0)+P(X=1)

=1-[4C0(12)4+4C1(12)(12)3]=1116

∴   n2p(x>1)=42×1116=11